Grade 12Algebra

Mathemitical step of
1/1^2 - 1/2^2 + 1/4^2 - 1/8^2......... 1/ infinity^2

Aatam
7 Years agoGrade 12

1 Answer

Aditya Gupta
7 Years ago
we see that this is just an infinite GP with common ratio r= – 1/2^2
so we apply the formula for sum of an infinte GP S= a/(1 – r)= 1/(1+1/4)= 4/5